With a Total Asset Turnover of 1.25X in 2018 for MEC Cor with no debt it generated sales of $15m. Its Basic Earning Power was 15%. With Corp. Taxes @35% Its ROE would be:

Select one:

a. 10.75%

b. 8.5%

c. 12.5%

d. 9.75%

e. None of the Above
Answer #1

Total Asset Turnover = Net Sales/Total Assets = 1.25
Total Assets = 15/1.25 = 12 m

BEP = EBIT/Total Assets = 0.15
EBIT =0.15*12 = 1.8 million

With no Debt , Equity = 12
EAT = 1.8*(1-35%) = 1.17

ROE = EAT/Equity = 1.17/12 = 9.75%

Option d is correct

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well